OpenStudy (anonymous):

Please Help me Find The greatest common monomial factor of the terms of the polynomial. 8x^3y^2 12x^2y^3 + 20x^2y^2 A. 4xy B. 4x^2y^2 C. 8x^2y^2 D.16xy

OpenStudy (cwrw238):

gcf of 8, 12 and 20 = 4 gcf of x^2 , x^3 and x^3 = x^2 gcf of y^2, y^3 and y^2 = ?

OpenStudy (cwrw238):

the gcf is the highest number / term which will divide into each of the numbers / terms.